Outline Introduction

Example Presentation Created with the Beamer Package

Till Tantau

April 17, 2008 Introduction

Overview of the Beamer Class

Outline Introduction

Overview of the Beamer Class

Features of the Beamer Class

▶ Normal LaTeX class. ▶ Easy overlays. ▶ No external programs needed.

Outline Introduction

Overview of the Beamer Class

Features of the Beamer Class

▶ Normal LaTeX class. ▶ Easy overlays. ▶ No external programs needed.

Outline Introduction

Overview of the Beamer Class

Features of the Beamer Class

▶ Normal LaTeX class. ▶ Easy overlays. ▶ No external programs needed.